OpenStudy (anonymous):

f(x)=-2x^3+6x^2-3 how do i find the x intercepts?

OpenStudy (anonymous):

I tried factoring it and it ends up -2x^2(x-3) and 3 isnt an x intercept.

OpenStudy (anonymous):

the x intercepts are a negative fraction and a positive fraction. Ive already graphed it im just trying to find out the long way.

OpenStudy (anonymous):

whhoops. -2(x-3)x^2

OpenStudy (anonymous):

\[f(x) = -2x^3+6x^2-3\] and find when \[f(x) = 0\]You can use the cubic formula for difficult cubics. (it's a bit too long to memorise) but I think that might be your only option! You can find it http://www.math.vanderbilt.edu/~schectex/courses/cubic/

OpenStudy (anonymous):

Oh lord. yea. i had it set equal to 0. i tried doing it 2x^3=6x^2-3 also as 3=-2x^2+6x^2

OpenStudy (anonymous):

neither seem to work.

OpenStudy (anonymous):

Other alternate forms are \[(6-2x)x^2-3\]and\[-2(x-2)^3-6(x-2)^2+5\]

OpenStudy (anonymous):

thank you :) Miscalculated.

OpenStudy (anonymous):

can you do an example of this equation? @_@

OpenStudy (anonymous):

Er, I'll give it a go. give me a minute

OpenStudy (anonymous):

Sure lol. thanks. Ive never done the cubic formula it looks like itll end up with a sqrt of something but kind of boggling my mind right now.

OpenStudy (anonymous):

Yeah, it's mega complicated, and can lead to complex numbers. (and does in your case too) however it should pop out with 3 numbers. With your formula\[a = -2\]\[b=6\]\[c=0\]\[d=-3\]Substituting that into crazy equation gives...(this might take a while)

OpenStudy (jack1):

... really? im getting x intercepts of: - 0.64178 0.83174 2.81

OpenStudy (anonymous):

wow, thats correct!

OpenStudy (anonymous):

i think hea found my y intercept but the x's were incorrect.

OpenStudy (jack1):

are we sure the equations right, that's a hell of a weird group of intercepts, usually they make these answers a bit cleaner

OpenStudy (anonymous):

yes. The original equation is f(x)=-2x^3+6x^2-3

OpenStudy (anonymous):

the first deriative is -6x^2+12x and the second is -12x+12

OpenStudy (anonymous):

my critical numbers were x=0 and x=2. min is -3 and max is at 5. seems to all work out?

OpenStudy (anonymous):

unless i derivative'd wrong. lol

OpenStudy (anonymous):

My math teacher is a jerk.

OpenStudy (anonymous):

I'm a little lost now, but basically \[-2x^3+6x^2-3=0\]The solutions are \[x=-0.642\]\[x=2.810\]Does that sort it?

OpenStudy (anonymous):

yea, but howd you get the x intercepts? D:

OpenStudy (anonymous):

the cubic?

OpenStudy (anonymous):

Yeah subbing the values from above into the mega long equation.

OpenStudy (jack1):

@hea you missed one hea, if you graph it: it passes through the x axis 3 times

OpenStudy (anonymous):

Thanks. This is a graphing question and im only in Calc 1. my teacher is sort of vague on these things but im assuming we are meant to graph it and find it from the graph then. its just weird because usually we do it the longhand way. the 0=-2x^3+6x^2-3 and solving for x. Do you guys think he meant for us to use the cubic formula?

OpenStudy (anonymous):

True my bad,\[x=0.832\]However I missed this because the solution is complex, however it is of course still valid

OpenStudy (anonymous):

Probably not, yeah I reckon teach meant for you to graph it and then estimate the result from that. Look in to various root finding methods, Newton's method, Bisection method and on and on. Is the kind of thing that your teacher has been going on about?

OpenStudy (anonymous):

nope. we had another question similar to this and he expects us to just estimate. thanks for all the help :)

OpenStudy (amistre64):

id get rid of the x^2 term, like cardano

OpenStudy (amistre64):

2x^3-6x^2+3 = 0 x^3-3x^2+3/2 = 0 ; let x = (u+1) (u+1)^3-(u+1)^2+3/2 = 0 u^3+u^2+u+1 -u^2-2u-1 +3/2 --------------- u^3 -u+3/2 = 0 u^3 = u - 3/2

OpenStudy (amistre64):

then there is a cubic formula that is derived .... similar to the quadratic formula